One cleaning fluid contains 6% ammonia, while a second contains 26% ammonia. In blending these fluids, how many liters of the second must be added to 10 liters of the first, if the ammonia in the mixture is to be 16%?
Answer by stanbon(75887)

One cleaning fluid contains 6% ammonia, while a second contains 26% ammonia. In blending these fluids, how many liters of the second must be added to 10 liters of the first, if the ammonia in the mixture is to be 16%?
-----------------
Equation:
active + active = active
0.06*10 + 0.26x = 0.16(10+x)
----
Multiply thru by 100 to get:
6*10 + 26x = 16*10 + 16x
10x = 10*10
x = 10 liters (amt. of 26% solution needed)
==============================================
Cheers,
Stan H.
=========